OpenStudy (anonymous):

Match Term Definition 128, 32, 8, 2, ... A) Geometric, common ratio is 3 1, 3, 9, 27, ... B) Geometric, common ratio is 0.25 5, 10, 15, 20, ... C) Arithmetic, common difference is -3 20, 17, 14, 11, ... D) Arithmetic, common difference is 5

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

ok for a geometric, you can take the 2nd term divided by the first term and that will be your common ratio, if all a_n+1 / a_n has the same common ratio that will tell you it is in fact a geometric sequence

OpenStudy (anonymous):

waits what is it @Nurali @billj5

OpenStudy (anonymous):

for an arithmetic, you can take the absolute value of 2nd term minus the first term and that will be your common difference, if all absolute value of a_n+1 - a_n has the same common difference that will tell you it is in fact an arithmetic sequence

OpenStudy (anonymous):

so what is the correct way @nurali @billj5

OpenStudy (anonymous):

what do u think @texaschic101

OpenStudy (texaschic101):

if the number is multiplied by a certain number to get to the next number, then it is an geometric sequence. If the number is added/subtracted to find the next number, it is an arithmetic sequence. 128,32,8,2 .. 32/128 = 0.25 check.. 128 * 0.25 = 32 32 * 0.25 = 8 the common ratio is 0.25 (geometric) 1,3,9,27 3/1 = 3 check.. 1 * 3 = 3 3 * 3 = 9 9 * 3 = 27 the common ratio is 3 (geometric) 5,10,15,20 10 - 5 = 5 15 - 10 = 5 20 - 15 = 5 common difference is 5 (arithmetic) 20,17,14,11 20 - 3 = 17 17 - 3 = 14 14 - 3 = 11 common difference is -3 (arithmetic)